# Point-PEFT: Parameter-Efficient Fine-Tuning for 3D Pre-trained Models

### v6

@@ 2 Related Work
The recent spotlight in the 3D domain has shifted from the supervised training [19, 18] towards self-supervised and large-model pre-training methods due to the challenge of data scarcity. These approaches adopt pretext tasks to pre-train large models, learning the latent representations embedded within point clouds. When fine-tuning on downstream tasks, such as classification [25, 24], segmentation [27], and 3D visual grounding [8], the pre-trained weights are optimized to acquire the knowledge related to the task. Some prior works [1, 26] utilize contrastive pretext tasks to pre-train the model, discriminating the different views of a single instance from views of other instances. Some concurrent works [33, 39] follow a training-free paradigm, leveraging pre-trained models like CLIP [21] for downstream tasks. More research works [17, 28] introduce the masked point modeling strategy for a stronger 3D encoder. Point-BERT [28] employs a point tokenizer to obtain the point tokens. Then the Encoder-Decoder architecture is used to model and predict masked point tokens. Point-MAE [17] and Point-M2AE [32] directly utilize Masked Autoencoders (MAE) [11], achieving superior representation capabilities. Recently, I2P-MAE [37], ACT [4], and Point-Bind [9] integrate rich knowledge from pre-trained multi-modal encoders to assist in 3D learning, indicating the potential of introducing external guidance. Joint-MAE [7] and PiMAE [3] conduct 2D-3D joint pre-training to incorporate cross-modal knowledge. Despite the success of 3D pre-training, the adaption for downstream tasks still demands the resource-intensive full fine-tuning method. Thus, we explore the PEFT techniques in the 3D domain for parameter-efficient fine-tuning.
@@ 3.3 Geometry-aware Adapter
Table 1: Real-world 3D Classification on ScanObjectNN. We report the number of learnable parameters (#Param) and the accuracy (%) on the ”PB-T50-RS” split of ScanObjectNN. ${}^{\dagger}$ indicates utilizing a stronger data augmentation [37] during fine-tuning.
@@ 4.2 Quantitative Analysis
As shown in Table 1, our Point-PEFT framework surpasses the full fine-tuning method with less than 5% trainable parameters. The improvements brought by our framework are +1.9%, +0.1%, and +1.0% for Point-BERT, ${}^{\dagger}$, and ${}^{\dagger}$ respectively, indicating our great advantages under complex 3D scenes by the extracted fine-grained geometric information. Compared to the full fine-tuning method, our Point-PEFT framework has the stronger ability to be adapted to the tasks related to the real-world scanned objects by pre-trained prior knowledge.
[…]
@@ 5 Ablation Study
In this section, we conduct extensive ablation studies to explore the effectiveness of different components within our Point-PEFT framework. We adopt ${}^{\dagger}$ as the pre-trained model, and report the classification accuracy (%) on the ”PB-T50-RS” split of the ScanObjectNN dataset.
@@
As shown in Table 4, to substantiate the effectiveness of each component, we conduct the ablation experiments by incrementally introducing components to the baseline (${}^{\dagger}$ model) until the complete Point-PEFT framework. The light gray row indicates the baseline with the transformer encoder, and the dark gray row represents the complete structure of the Point-PEFT framework. Introducing either the Point-prior Prompt or the Geometry-aware Adapter component leads to a certain degree of performance degradation compared to full fine-tuning. When both components are utilized, the performance has been improved to 88.6% with a gain of +0.5%. Further adding Bias Tuning can result in an additional +0.5% improvement. Therefore, the experiments indicated the effectiveness of each design within the Point-PEFT framework.
@@ Appendix B Additional Experiments
[…]
@@
[…]
@@ Appendix C Additional Ablation Study
We further conduct ablation study on the positions and depth of the proposed Geometry-aware Adapter in Table 9, where we utilize the final version of Point-PEFT as the baseline and compare the accuracy(%) on the ”PB-T50-RS” split of the ScanObjectNN with ${}^{\dagger}$ as the pre-trained model. As shown in Table 9, the last position within the block and the depth of 12 perform the best. That is because, compared to other positions, after the global interactions among point cloud features through pre-trained self-attention layer and Feed-forward Networks(FFN), the local interactions within the adapter could further boost the extraction of the fine-grained geometric information, which facilitates the fine-tuning process. And the deeper insertion could utilize the high-level knowledge to grasp the better understanding of 3D structures. Note that, for all pre-trained point cloud models, we insert the Geometry-aware Adapter in the last position across all blocks.

### v7

@@ 2 Related Work
The recent spotlight in the 3D domain has shifted from the supervised training [19, 18] towards self-supervised and large-model pre-training methods due to the challenge of data scarcity. These approaches adopt pretext tasks to pre-train large models, learning the latent representations embedded within point clouds. When fine-tuning on downstream tasks [20, 21], such as classification [27, 26], segmentation [29], and 3D visual grounding [8], the pre-trained weights are optimized to acquire the knowledge related to the task. Some prior works [1, 28] utilize contrastive pretext tasks to pre-train the model, discriminating the different views of a single instance from views of other instances. Some concurrent works [35, 41] follow a training-free paradigm, leveraging pre-trained models like CLIP [23] for downstream tasks. More research works [17, 30] introduce the masked point modeling strategy for a stronger 3D encoder. Point-BERT [30] employs a point tokenizer to obtain the point tokens. Then the Encoder-Decoder architecture is used to model and predict masked point tokens. Point-MAE [17] and Point-M2AE [34] directly utilize Masked Autoencoders (MAE) [11], achieving superior representation capabilities. Recently, I2P-MAE [39], ACT [4], and Point-Bind [9] integrate rich knowledge from pre-trained multi-modal encoders to assist in 3D learning, indicating the potential of introducing external guidance. Joint-MAE [7] and PiMAE [3] conduct 2D-3D joint pre-training to incorporate cross-modal knowledge. Despite the success of 3D pre-training, the adaption for downstream tasks still demands the resource-intensive full fine-tuning method. Thus, we explore the PEFT techniques in the 3D domain for parameter-efficient fine-tuning.
@@ 3.3 Geometry-aware Adapter
Table 1: Real-world 3D Classification on ScanObjectNN. We report the number of learnable parameters (#Param) and the accuracy (%) on the ”PB-T50-RS” split of ScanObjectNN. indicates utilizing a stronger data augmentation [39] during fine-tuning.
@@ 4.2 Quantitative Analysis
As shown in Table 1, our Point-PEFT framework surpasses the full fine-tuning method with less than 5% trainable parameters. The improvements brought by our framework are +1.9%, +0.1%, and +1.0% for Point-BERT, Point-M2AE, and Point-MAE respectively, indicating our great advantages under complex 3D scenes by the extracted fine-grained geometric information. Compared to the full fine-tuning method, our Point-PEFT framework has the stronger ability to be adapted to the tasks related to the real-world scanned objects by pre-trained prior knowledge.
Local Aggregation Self-Attn. MLPfinal Acc. (%)
@@ 5 Ablation Study
In this section, we conduct extensive ablation studies to explore the effectiveness of different components within our Point-PEFT framework. We adopt Point-MAE as the pre-trained model, and report the classification accuracy (%) on the ”PB-T50-RS” split of the ScanObjectNN dataset.
@@
As shown in Table 4, to substantiate the effectiveness of each component, we conduct the ablation experiments by incrementally introducing components to the baseline (Point-MAE model) until the complete Point-PEFT framework. The light gray row indicates the baseline with the transformer encoder, and the dark gray row represents the complete structure of the Point-PEFT framework. Introducing either the Point-prior Prompt or the Geometry-aware Adapter component leads to a certain degree of performance degradation compared to full fine-tuning. When both components are utilized, the performance has been improved to 88.6% with a gain of +0.5%. Further adding Bias Tuning can result in an additional +0.5% improvement. Therefore, the experiments indicated the effectiveness of each design within the Point-PEFT framework.
@@ Appendix B Additional Experiments
Table 8: Part Segmentation on ShapeNetPart [29]. ‘mIoUC’ (%) and ‘mIoUI’ (%) denote the mean IoU across all part categories and all instances in the dataset, respectively. denotes that we do not include classification heads into the parameter calculation.
@@
To evaluate the understanding capacity of Point-PEFT on dense 3D data, we further conduct part segmentation experiments on ShapeNetPart [29] that contains 14,007 and 2,874 samples for training and validation, respectively. We follow previous works to adopt the same segmentation head and sample 2048 points for each object as input. For each pre-trained model, we fine-tune Point-PEFT for 300 epochs with a batch size of 16, and adopt the learning rate as 2$\times$10-4. In Table 8, we report the mean IoU across all part categories and instances, denoted as mIoUC and mIoUI, respectively. As reported, on each pre-trained model, our Point-PEFT attains the competitive performance on both metrics, which illustrates the strong understanding capacity of our designs on the fine-grained dense data. Compared to the full fine-tuning method, the number of learnable parameters has dropped 98% and 97% for the Point-MAE and Point-M2AE respectively.
@@ Appendix C Additional Ablation Study
We further conduct ablation study on the positions and depth of the proposed Geometry-aware Adapter in Table 9, where we utilize the final version of Point-PEFT as the baseline and compare the accuracy(%) on the ”PB-T50-RS” split of the ScanObjectNN with Point-MAE as the pre-trained model. As shown in Table 9, the last position within the block and the depth of 12 perform the best. That is because, compared to other positions, after the global interactions among point cloud features through pre-trained self-attention layer and Feed-forward Networks(FFN), the local interactions within the adapter could further boost the extraction of the fine-grained geometric information, which facilitates the fine-tuning process. And the deeper insertion could utilize the high-level knowledge to grasp the better understanding of 3D structures. Note that, for all pre-trained point cloud models, we insert the Geometry-aware Adapter in the last position across all blocks.